# Non-Linearities in International Spillovers of the ECB’s Monetary Policy

### v1

@@ 3.1 The Structural Model
We postulate that the data generating process, for each country, is a structural vector auto-regressive model of the following form (similar to the one studied in Gonçalves et al. [2021]):
$A_{0}w_{t}=a+A(L)w_{t-1}+C(L)f(x_{t})+\eta_{t}$ (1)
Where $w_{t}=\langle x_{t},y_{t},z_{t}\rangle$, $x_{t}$ is a vector of identified shocks, $y_{t}$ is a vector of country specific endogenous variables and $z_{t}$ are the euro-area endogenous variables.
@@
The elements of $y_{t}$ and $z_{t}$ do not impact $x_{t}$ contemporaneously but $x_{t}$ does affects $y_{t}$ and $z_{t}$ contemporaneously. We do not restrict the contemporaneous relationship between $y_{t}$ and $z_{t}$ as we only require a block-recursive DGP to identify the impulse response to a shock in $x_{t}$.
@@
$\eta_{t}=\langle u_{t},\epsilon_{t},v_{t}\rangle\sim i.i.d\left(\begin{bmatrix}0\\ 0\\ 0\end{bmatrix},\begin{bmatrix}\mathbf{I}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\Sigma_{2,2}&\Sigma_{2,3}\\ \mathbf{0}&\Sigma_{3,2}&\Sigma_{3,3}\\ \end{bmatrix}\right)$ (5)
@@ 3.2 The Linear Local Projection and Model Selection
We estimate the impulse responses of interest using the local projection method developed by Dufour and Renault [1998] and popularized by Jordà [2005]. In the first baseline specification we treat $C_{2}(L)$ as a zero matrix, completely ignoring non-linearities. We refer to this as the linear specification which yields the following local projection.

### v2

@@ 3.1 The Structural Model
We postulate that the data generating process, for each country j, is a structural vector auto-regressive model of the following form (similar to the one studied in Gonçalves et al. [2021] and the VARX of Dees et al. [2007]):
$A_{0}w_{j,t}=a+A(L)w_{j,t-1}+C(L)f(x_{t})+\eta_{j,t}$ (1)
Where $w_{j,t}=\langle x_{t},y_{j,t},z_{t}\rangle$, $x_{t}$ is a vector of identified shocks, $y_{j,t}$ is a vector of country specific endogenous variables and $z_{t}$ are the euro-area endogenous variables.
@@
The elements of $y_{j,t}$ and $z_{t}$ do not impact $x_{t}$ contemporaneously but $x_{t}$ does affects $y_{j,t}$ and $z_{t}$ contemporaneously. We do not restrict the contemporaneous relationship between $y_{j,t}$ and $z_{t}$ as we only require a block-recursive DGP to identify the impulse response to a shock in $x_{t}$.
@@
$\eta_{j,t}=\langle u_{t},\epsilon_{j,t},v_{t}\rangle\sim i.i.d\left(\begin{bmatrix}0\\ 0\\ 0\end{bmatrix},\begin{bmatrix}\mathbf{I}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&\Sigma_{2,2}&\Sigma_{2,3}\\ \mathbf{0}&\Sigma_{3,2}&\Sigma_{3,3}\\ \end{bmatrix}\right)$ (5)
@@ 3.2 The Linear Local Projection and Model Selection
We estimate the impulse responses of interest using the method of local projections (Dufour and Renault [1998] and Jordà [2005]). In the first baseline specification we treat $C_{2}(L)$ as a zero matrix, completely ignoring non-linearities. We refer to this as the linear specification which yields the following local projection.